OpenStudy (anonymous):

cube root of 56 x to the fourteenth power

OpenStudy (anonymous):

the cube root can be combined with the exponent as a fraction cube root of x to the fourteenth power can be written as x^14/3 so cube root of 56x^14 (56x)^14/3

OpenStudy (anonymous):

\[\huge (\sqrt[3]{56x})^{14}\] \[\huge (56x)^{\frac{14}{3}}\] \[\huge (56)^{\frac{14}{3}}*(x)^{\frac{14}{3}}\] \[\huge (28*2)^{\frac{14}{3}}*(x)^{\frac{14}{3}}\] \[\huge (14*2*2)^{\frac{14}{3}}*\sqrt[3]{x^{14 }}\] \[\huge (7*2^3 )^{\frac{14}{3}}*\sqrt[3]{x^{12}*x^2 }\] \[\huge (7)^{\frac{14}{3}}*(2^3 )^{\frac{14}{3}}*\sqrt[3]{x^{12}} \sqrt[3]{x^2 }\] \[\huge \sqrt[3]{7^{14}}*(2^{14}) *x^4 \sqrt[3]{x^2 }\]

OpenStudy (anonymous):

Thank you so much!!!